OpenStudy (anonymous):

Harriet Amato operates a coffee shop. One of her customers wants to buy two kinds of beans, Arabian Mocha, and Columbian Decaf. If she wants twice as much mocha as columbian, how much of each can she buy for a total of $87.50?

OpenStudy (anonymous):

x+y = 87.50 and 2x = y use that to solve for each.

OpenStudy (anonymous):

That's sort of confusing... I've tried using X = Columbian, 2X = Arab and I set it up as: X + 2x = 87.50 But it's not coming out right...Could you please explain a little bit more?

OpenStudy (anonymous):

x = col y = a the total that can be spent on both is 87.50 so x + y = 87.50, but you also want 2 times more of one than the other so 2x = y, so you substitute 2x in for y in the first equation and solve. Follow?

OpenStudy (anonymous):

How come I end up with 29.16 after I solve it? I set it up as x+2x=87.50 then combined the terms, and then divided. Am I doing it wrong? The book says the answer is Arab=7lb, and Columbian=3.5lb ??

OpenStudy (anonymous):

You are good so far

OpenStudy (anonymous):

So it is 29.16$ for both of them

OpenStudy (anonymous):

How do I divide that up then?

OpenStudy (anonymous):

Sorry I mean 29.16$ for Col

OpenStudy (anonymous):

And for Mocha u need to multiply with 2 which gives U $58.32

OpenStudy (anonymous):

How do I get the # of lb's out of the prices though? This problem is confusing.

OpenStudy (anonymous):

Seems like Incomplete question for me